OpenStudy (anonymous):

The Substitution Method - Solving Systems of Equations by Graphing: y = 2x - 5 3y - x = 5

OpenStudy (rulnick):

By substitution, put 2x-5 in for y in the second equation: 3(2x-5) - x = 5 solve to get 6x-15-x=5 5x=20 x=4 back-substitute to find y: y=2(4)-5 y=3 So the point of intersection (solution point) is (x,y)=(4,3).

OpenStudy (anonymous):

Wow, Thanks!

OpenStudy (rulnick):

To graph, the first equation has slope 2 and y-intercept -5, the second has slope 1/3 and intercept 5.

OpenStudy (anonymous):

y = 2x - 5 3y - x = 5 -> 3( 2x -5) -x = 5 -> 6x - 15 -x = 5 -> 5x = 20 => x = 20/5 = 4 y = 2 * 4 - 5 = 3 Thus intersection (4, 3)